OpenStudy (anonymous):

show how to solve: √x+2+4=x

OpenStudy (anonymous):

\[\sqrt{x+2}+4=x\]

OpenStudy (anonymous):

is this your question??

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

id try to get the sqrt part all by itself then undo it by squaring both sides

OpenStudy (anonymous):

@TiaW

OpenStudy (anonymous):

yes??

OpenStudy (amistre64):

squaring the equation does introduce a "false" result which would have to be determined at the end of the setup

OpenStudy (anonymous):

(√x+2)(√4)

OpenStudy (anonymous):

@03453660 yes?

OpenStudy (anonymous):

ok coming to you

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

wait.. waat?

OpenStudy (anonymous):

i know its 7. but i dont know how to get it

OpenStudy (amistre64):

√(x+2) + 4 = x to get the sqrt all alone, we would have to subtract off 4 right?

OpenStudy (anonymous):

yeah

OpenStudy (amistre64):

√(x+2) + 4 = x - 4 -4 --------------- √(x+2) = x -4 to undo a sqrt, we need to square it

OpenStudy (amistre64):

[√(x+2)]^2 = (x -4)^2 x+2 = (x - 4)^2 can you expand the right side of this?

OpenStudy (anonymous):

nope. im lost

OpenStudy (amistre64):

at what point did you get lost at?

OpenStudy (anonymous):

never mind. i was lookin at it wrong.. so it'll be: x+2 = (x - 4)^2 x+2=x^2+16 x=x^2+14 x-x2=14??

OpenStudy (amistre64):

something like that yes, but i think there are a few missteps along the way x+2 = (x - 4)^2 x+2 = x^2 - 8x +16 -x - x ------------------- 2 = x^2 - 9x +16 -2 -2 ----------------- 0 = x^2 -9x +14

OpenStudy (amistre64):

there are different ways that we can solve a quadratic; factoring might be a good method for this. 2 numbers that add to -9 and multiply to 14 ; -7 and -2 (x-7)(x-2) = 0 ; when x=7 or x=2

OpenStudy (anonymous):

where 8x come from?

OpenStudy (amistre64):

the most used method for this is the "foil" method; which is a memory aid for firsts lasts outsides insides